9 Graham Potato Company has projected sales of $7,200 in September, $11,000 in October, $17,200 in November, and $13,200 in December. Of the companys sales, 20 percent are paid for by cash and 80 percent are sold on credit. Experience shows that 40 percent of accounts receivable are paid in the month after the sale, while the remaining 60 percent are paid two months after. Determine collections for November and December. 9.09 points Also assume Grahams cash payments for November and December are $14,500 and $7,000, respectively. The beginning cash balance in November is $5,000, which is the desired minimum balance. a. Prepare a cash receipts schedule for November and December eBook Graham Potato Company Cash Receipts Schedule October Hint September November December Sales 7,200 S 11,000 $ 17,200 S 13,200 10,560 2,640 5,504 5,280 13,424 Print Credit sales 5,760 8,800 13,760 Cash sales One month after sale Two months after sale 3,440 S 3,520 3,456 Total cash receipts 10,416$
Answer #1

Credit sales

Credit Sales 80% of Sales

September =7200*80%=$5760

October =11000*80%=$8800

November =17200*80%=$13760

December=13200*80%=$10560

Cash sales

Credit Sales 20% of Sales

November =17200*20%=$3440

December=13200*20%=$2640

One Month After Sales

40% of credit sales are paid after one month

November=$8800*40%=$3520

December =$13760*40%=$5504

Two Month After Sales

November=$5760*60%=$3456

December =$8800*60%=$5280

Total Cash Recripts

Cash Sales +One Month After Sales+Two Month After Sales

November=3440+3520+3456=$10416

December=2640+5504+5280=$13424

SfBX6DWUp34AAAAASUVORK5CYII=